# Super-LIO: A Robust and Efficient LiDAR-Inertial Odometry System with a Compact Mapping Strategy

### v1

@@ IV-A Experimental Setup
m2s1 0.38 0.40 0.38 0.32 0.32 0.33 0.75
m2s2 2.80 3.05 2.86 2.89 3.11 2.97 1.48
m2s3 0.14 0.15 0.20 0.20 0.17 0.20 0.42
m2s4 0.64 0.67 0.45 0.47 0.60 0.43 0.84
m2s5 0.38 0.39 0.38 0.37 0.39 0.27 0.42
m2h1 0.29 0.29 0.30 0.30 0.30 0.31 0.21
m2h5 1.16 1.17 1.20 1.20 1.19 1.17 0.29
m2d1 0.46 0.45 0.47 0.47 0.45 0.44 0.14
m2d2 0.32 0.34 0.36 0.36 0.34 0.37 0.06
nclt1 1.69 1.76 1.72 1.73 2.02 2.48 7.58
nclt2 1.30 1.32 1.42 1.48 1.32 1.17 3.19
nclt3 1.82 1.93 2.29 2.33 2.42 1.88 6.12
mcd0 0.54 0.63 0.48 0.57 0.57 1.29 3.20
mcd2 0.72 0.79 1.08 1.41 0.95 1.69 1.78
mcd3 0.50 0.56 0.58 0.75 0.56 0.92 1.46
mcd4 0.60 0.63 0.96 1.12 0.70 0.75 2.42
eee1 0.12 0.12 0.08 0.08 0.16 0.26
nya1 0.07 0.09 0.06 0.06 0.13 0.07 0.19
sbs1 0.09 0.09 0.09 0.09 0.13 0.09 0.22
Avg 0.74 0.78 0.81 0.85 0.83 0.93 1.63
@@
IV-B Accuracy Evaluation
Accuracy is evaluated on public datasets with ground-truth trajectories, as introduced in Section IV-A. We use the Root Mean Square Error (RMSE) between estimated and ground-truth poses, computed by the Evo toolbox [38], as the evaluation metric.
For brevity, dataset names are simplified in Table I: for example, M2DGR sequences are denoted by the prefix “m2” and the NTU prefix is omitted in the last three rows. The same naming rule applies in the following analysis.
The evaluated datasets cover a wide spectrum of LiDAR types, platforms, and scenarios. M2DGR [34] and NCLT [35] both employ a Velodyne HDL-32E spinning LiDAR. M2DGR consists of low-speed ground-robot runs, where we evaluate outdoor street and indoor corridor scenes. NCLT provides long-duration trajectories (up to 110 minutes and 7.58 km), allowing us to assess accuracy under extended continuous operation. The MCD [36] dataset uses a Livox MID-70 solid-state LiDAR with a small FoV, recorded on a high-speed terrain vehicle; the tested sequences exceed 1 km, with peak speed $\sim$10 m/s and angular velocity $\sim$2.9 rad/s, stressing both robustness and accuracy under fast motion. Finally, the NTU [37] dataset is collected with a UAV equipped with an Ouster-16 LiDAR, yielding relatively sparse aerial point clouds; this setting is particularly challenging, and iG-LIO [18] fails on one sequence (eee1).
Table I reports the RMSE results across multiple sequences. Super-LIO achieves the best average accuracy among all tested methods, despite relying on a significantly more compact map representation. In several sequences, it even surpasses state-of-the-art baselines, demonstrating the effectiveness of the OctVox map. The ablation variant Super-LIO* shows a modest yet expected degradation, confirming that the proposed HKNN search improves the data association quality during scan-to-map alignment. By contrast, FAST-LIO2* exhibits a slight accuracy drop compared with FAST-LIO2 [14], which is mainly attributed to more aggressive multi-threading that introduces small numerical inconsistencies in scan-to-map optimization. While accuracy is important, the focus of this work is on optimizing efficiency, as discussed in the following sections.
@@ IV-C Efficiency Analysis
To evaluate runtime efficiency, we measure both per-frame processing time and CPU usage on the X86 laptop and the embedded NVIDIA Orin NX. In addition to public datasets, we introduce ten representative self-collected sequences, denoted with the prefix “se” for efficiency and robustness evaluation. These sequences are captured with a Livox MID-360 LiDAR, which employs a non-repetitive scanning pattern. Due to its lightweight design, this sensor is well suited for small-scale platforms such as UAVs, making the collected datasets more practically relevant.
Figure 1 provides an overview of average processing time and CPU utilization across methods, while Table II reports detailed results for each sequence, including per-frame processing time and standard deviation on both X86 and ARM platforms. From these results, Super-LIO achieves a clear efficiency advantage: it consumes fewer CPU resources than FAST-LIO2 on both platforms, while delivering $3.7\times$ and $4.2\times$ speedups on X86 and ARM, respectively.
TABLE II: Per-frame Runtime Statistics on x86 (AMD 5800H, 5× playback) and ARM (Cortex-A78AE), in milliseconds
Seq Super-LIO Super-LIO* FAST-LIO2 FAST-LIO2* Faster-LIO iG-LIO
x86 ARM x86 ARM x86 ARM x86 ARM x86 ARM x86 ARM
Avg Std Avg Std Avg Std Avg Std Avg Std Avg Std Avg Std Avg Std Avg Std Avg Std Avg Std Avg Std
m2s1 4.94 1.31 19.18 2.33 5.85 1.63 21.46 2.67 19.72 3.78 93.28 7.43 8.84 1.63 30.51 2.62 14.55 2.71 45.58 6.39 7.37 2.60 20.89 3.51
m2s2 4.38 1.23 16.97 3.19 5.01 1.52 18.42 2.94 17.77 3.52 77.26 12.98 7.83 1.76 27.61 3.96 12.77 2.48 37.07 4.26 7.20 2.46 21.45 3.60
m2s3 5.20 1.19 20.01 2.93 6.21 1.42 21.17 2.91 24.66 2.53 92.78 9.05 9.46 1.21 30.57 2.84 18.33 3.65 44.39 6.29 8.25 2.15
m2s4 4.75 1.54 20.35 2.66 5.28 1.55 21.95 2.67 19.68 4.10 93.85 5.25 8.26 1.66 30.48 2.74 13.27 3.27 42.62 4.24 6.95 2.61 21.57 4.49
m2s5 4.46 1.40 15.36 4.02 4.67 1.38 16.22 4.31 17.23 3.89 66.52 12.87 7.87 1.88 25.99 3.39 11.19 2.21 34.17 4.24 6.20 1.92 19.92 3.32
m2h1 2.85 0.99 10.16 2.69 3.23 1.09 11.25 2.90 11.89 1.65 45.72 4.39 6.81 1.90 22.77 2.81 12.15 2.68 30.26 4.06 5.52 1.37
m2h5 2.55 0.88 9.58 2.63 3.62 1.05 11.16 3.03 10.65 1.84 43.49 3.85 6.37 1.75 22.38 2.73 11.53 2.62 29.01 3.66 4.89 1.40
m2d1 2.81 1.05 11.69 3.62 3.23 1.07 12.47 3.90 11.24 3.04 50.12 11.03 6.43 1.88 22.90 3.76 10.31 1.89 29.46 4.55 4.68 1.48
m2d2 2.97 1.10 10.74 3.28 3.35 1.15 11.69 3.88 10.81 2.16 46.20 10.04 6.51 1.97 22.12 3.37 8.26 1.67 28.47 3.89 4.72 1.40 15.03 3.54
nclt1 4.25 1.60 7.97 3.02 5.30 1.91 9.81 3.63 13.27 4.64 31.01 8.34 7.29 2.26 14.09 4.07 9.82 2.95 15.45 4.53 5.92 2.98
nclt2 4.20 1.51 9.06 3.13 4.99 1.79 10.84 3.97 13.85 3.01 34.86 9.67 7.48 2.21 15.64 4.46 9.96 2.44 17.93 5.07 6.02 2.32
nclt3 4.14 1.60 10.88 2.95 5.15 1.64 12.38 3.26 13.48 37.15 42.75 8.11 7.19 2.19 19.97 3.91 9.92 3.06 23.98 4.53 5.37 2.15
mcd0 1.59 1.00 6.07 1.62 1.97 1.34 8.68 2.06 5.37 1.26 20.60 4.69 3.06 1.17 10.38 3.09 3.12 1.36 10.55 2.97 2.98 1.47
mcd2 1.84 0.92 7.06 2.27 2.31 0.94 7.30 1.97 5.59 1.52 24.66 5.64 3.09 1.16 11.24 2.82 3.37 1.94 12.40 4.07 3.05 1.84
mcd3 1.84 0.89 7.22 1.63 2.10 0.93 7.61 2.47 5.51 1.39 23.80 5.11 3.03 1.17 11.05 2.91 3.14 1.13 12.38 3.32 3.17 1.85
mcd4 1.65 0.77 5.86 2.25 1.95 0.89 6.83 3.24 5.52 1.53 22.04 6.11 3.11 1.22 10.87 3.25 3.16 1.59 11.26 3.59 3.15 7.43
eee1 3.22 1.11 9.91 2.73 3.68 1.24 10.16 4.06 10.54 2.33 33.84 11.84 5.47 1.59 14.37 4.47 9.38 2.65 17.79 7.09
nya1 2.44 0.96 7.15 1.94 2.69 1.10 8.84 2.59 8.00 1.75 23.19 6.33 4.85 1.56 12.16 3.72 8.20 2.63 13.64 4.73 3.66 1.16
sbs1 2.74 1.06 8.77 2.67 3.08 1.10 10.53 3.90 8.32 1.84 30.16 8.73 4.84 1.51 12.93 4.22 7.39 1.88 16.32 6.52 3.78 1.30
se01 3.20 1.23 13.69 3.16 3.30 1.15 16.09 3.35 11.10 1.61 52.33 6.50 5.24 1.19 19.64 4.08 7.14 1.30 26.33 3.54 4.84 1.75
se02 2.76 1.00 9.99 2.98 2.87 1.03 13.17 3.27 9.42 1.88 40.56 5.91 4.91 1.33 15.56 3.46 6.50 1.53 21.94 3.31 3.96 1.42
se03 2.53 0.94 9.14 2.10 2.68 0.96 12.77 2.73 8.74 1.42 38.43 3.45 5.13 1.46 15.57 2.52 6.14 1.50 21.11 2.71 3.85 1.44
se04 0.76 0.33 3.81 1.91 0.80 0.32 4.72 2.47 2.99 0.91 15.89 6.40 1.73 0.68 6.62 2.78 2.54 0.87 8.65 2.84 1.67 0.63
se05 1.07 0.47 4.79 2.14 1.12 0.48 5.72 2.61 4.50 1.53 19.91 6.45 2.62 1.07 9.20 2.85 3.12 1.20 11.67 3.16 2.19 0.92 8.21 2.90
se06 0.75 0.36 3.22 1.47 0.74 0.34 4.45 1.70 3.09 1.30 14.32 6.69 1.97 0.87 6.79 2.67 2.35 1.12 8.47 3.36 1.67 0.71
se07 1.16 0.46 4.99 1.35 1.18 0.46 5.92 1.77 5.43 1.42 25.36 4.62 3.63 1.44 12.17 3.22 4.23 1.65 14.35 3.00 2.61 1.00
se08 0.69 0.45 2.19 1.19 0.72 0.46 2.83 1.58 3.26 2.10 9.06 3.65 2.07 1.27 5.65 2.48 2.74 1.43 7.34 2.63
se09 0.69 0.27 3.54 1.83 0.74 0.26 4.16 2.08 3.07 0.91 15.54 4.56 1.88 0.63 6.84 1.84 2.41 0.83 8.93 2.31 1.56 0.57
se10 0.64 0.33 3.25 1.48 0.63 0.31 3.29 1.33 3.01 1.64 14.60 5.93 2.18 1.23 7.32 2.74 2.59 1.38 9.56 3.41
AVG 2.66 0.96 9.40 2.45 3.05 1.05 10.75 2.87 9.92 3.37 39.38 7.09 5.14 1.48 16.32 3.23 7.57 1.99 21.07 4.08 4.43 1.86 17.85 3.56
RATE 1x 1x 1x 1x 1.2x 1.1x 1.1x 1.2x 3.7x 3.5x 4.2x 2.9x 1.9x 1.5x 1.7x 1.3x 2.9x 2.1x 2.2x 1.7x 1.8x 1.9x 1.9x 1.6x
It is worth noting that many prior works focus primarily on reducing average processing time, while methods such as Faster-LIO [14] rely on aggressive concurrency to maintain real-time performance. However, such strategies often impose a heavy burden on CPU resources, especially on embedded platforms. To enable fairer evaluation of overall efficiency, we further introduce a composite metric, relative efficiency, defined as follows:
$\eta=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{t_{i}\cdot u_{i}}$
where $t_{i}$ denotes the per-frame processing time measured in milliseconds and $u_{i}$ represents the normalized CPU utilization (011), balancing frame rate against computational cost. Based on this definition, Table III summarizes the relative efficiency of all methods on both the X86 and ARM platforms. Super-LIO delivers $2.4\times$ and $1.61\times$ the relative efficiency of iG-LIO on X86 and ARM, respectively, demonstrating its consistent advantage across platforms.
TABLE III: Relative efficiency $\uparrow$ on X86 and ARM platforms.
@@
X86 1.99 1.56 0.42 0.66 0.30 0.83
ARM 1.71 1.37 0.39 0.66 0.40 1.06
@@ IV-D Robustness Analysis
We also rigorously evaluated performance across both X86 and ARM platforms, confirming the cross-platform consistency of Super-LIO, as shown in Table II. In contrast, although iG-LIO delivers competitive real-time performance and accuracy on X86, it frequently fails on ARM.

### v2

@@ IV-B Accuracy Evaluation
m2s1 0.384 0.401 0.381 0.325 0.323 0.328 0.752
m2s2 2.801 3.052 2.862 2.891 3.109 2.968 1.483
m2s3 0.139 0.148 0.198 0.195 0.167 0.201 0.424
m2s4 0.641 0.671 0.455 0.470 0.601 0.429 0.840
m2s5 0.382 0.394 0.379 0.369 0.394 0.274 0.421
m2h1 0.291 0.294 0.303 0.302 0.298 0.305 0.214
m2h5 1.162 1.166 1.198 1.197 1.194 1.170 0.285
m2d1 0.456 0.454 0.465 0.465 0.450 0.442 0.143
m2d2 0.321 0.339 0.364 0.364 0.341 0.371 0.057
nclt1 1.692 1.763 1.719 1.730 2.022 2.475 7.582
nclt2 1.296 1.316 1.419 1.481 1.325 1.170 3.186
nclt3 1.821 1.934 2.289 2.330 2.421 1.876 6.121
mcd0 0.541 0.634 0.484 0.572 0.574 1.289 3.197
mcd2 0.721 0.790 1.085 1.413 0.951 1.692 1.784
mcd3 0.498 0.560 0.583 0.748 0.557 0.921 1.459
mcd4 0.604 0.629 0.964 1.117 0.695 0.752 2.421
eee1 0.119 0.123 0.079 0.083 0.164 0.261
nya1 0.069 0.078 0.063 0.064 0.127 0.072 0.193
sbs1 0.086 0.086 0.087 0.088 0.131 0.088 0.219
Avg 0.738 0.781 0.809 0.853 0.834 0.934 1.634
@@
Accuracy is evaluated on public datasets with ground-truth trajectories, as introduced in Section IV-A. We use the Root Mean Square Error (RMSE) between estimated and ground-truth poses, computed by the Evo toolbox [evo], as the evaluation metric.
For brevity, dataset names are simplified in Table I: for example, M2DGR [m2dgr] sequences are denoted by the prefix “m2” and the NTU VIRAL [NTU] prefix is omitted in the last three rows. The same naming rule applies in the following analysis.
The evaluated datasets cover a wide spectrum of LiDAR types, platforms, and scenarios. M2DGR [m2dgr] and NCLT [nclt] both employ a Velodyne HDL-32E spinning LiDAR. M2DGR consists of low-speed ground-robot runs, where we evaluate outdoor street and indoor corridor scenes. NCLT provides long-duration trajectories (up to 110 minutes and 7.58 km), allowing us to assess accuracy under extended continuous operation. The MCD [MCD] dataset uses a Livox MID-70 solid-state LiDAR with a small FoV, recorded on a high-speed terrain vehicle; the tested sequences exceed 1 km, with peak speed $\sim$10 m/s and angular velocity $\sim$2.9 rad/s, stressing both robustness and accuracy under fast motion. Finally, the NTU VIRAL [NTU] dataset is collected with a UAV equipped with an Ouster-16 LiDAR, yielding relatively sparse aerial point clouds; this setting is particularly challenging, and iG-LIO [ig_lio] fails on one sequence (eee1).
Table I reports the RMSE results across multiple sequences. Super-LIO achieves the best average accuracy among all tested methods, despite relying on a significantly more compact map representation. In several sequences, it even surpasses state-of-the-art baselines, illustrating that the OctVox map provides a reliable geometric representation. The ablation variant Super-LIO* shows a modest yet expected degradation, confirming that the proposed HKNN search improves the data association quality during scan-to-map alignment. By contrast, FAST-LIO2* exhibits a slight accuracy drop compared with FAST-LIO2 [fast_lio2], which is mainly attributed to more aggressive multi-threading that introduces small numerical inconsistencies in scan-to-map optimization.
@@ IV-C Efficiency Analysis
This section evaluates the runtime performance and system-resource usage of all methods on both the X86 laptop and the embedded NVIDIA Orin NX. Experiments cover all public datasets as well as ten additional self-collected (prefix “se”) Livox-MID-360 sequences.
Table II and Table III report the scene-level average per-frame processing time on the x86 laptop and the NVIDIA Orin NX, respectively. For each scene, all available sequences are averaged to obtain a single representative runtime measurement. Across all public and self-collected scenes, Super-LIO achieves the lowest processing latency on both platforms, demonstrating a consistent efficiency gain across hardware architectures. As illustrated in Fig. 1, it delivers $3.7\times$ and $4.2\times$ speedups over FAST-LIO2 on X86 and ARM, respectively, while also requiring fewer CPU resources. In contrast, iG-LIO performs competitively on x86 but incurs higher latency on ARM.
TABLE II: X86 per-frame runtime(ms), averaged across sequence groups.
@@
m2s 4.75 5.40 19.41 8.85 14.82 7.79
m2h 2.70 3.43 11.27 6.59 11.84 5.21
m2d 2.89 3.29 11.02 6.47 9.29 4.70
nclt 4.19 5.14 13.53 7.32 9.90 5.77
mcd 1.73 2.08 5.50 3.06 3.16 3.23
ntu 2.80 3.15 8.95 4.39 8.32 3.72
se 1.88 1.97 6.59 3.01 4.24 3.79
Avg 2.99 3.49 10.90 5.67 8.80 4.89
TABLE III: ARM per-frame runtime(ms), averaged across sequence groups.
Super-LIO Super-LIO* FAST-LIO2 FAST-LIO2* Faster-LIO iG-LIO
m2s 18.77 19.84 84.74 28.63 38.81 40.47
m2h 9.87 11.21 44.61 22.58 29.63 22.45
m2d 11.22 12.08 48.16 22.51 28.97 24.60
nclt 9.30 11.01 36.87 16.57 19.79 17.02
mcd 6.03 7.35 20.40 10.57 10.65 10.01
ntu 8.94 9.84 29.06 12.82 15.25
se 7.17 8.29 25.52 9.64 12.64 14.89
Avg 10.47 11.66 41.34 17.76 22.53 21.24
TABLE IV: Relative efficiency $\uparrow$ on X86 and ARM platforms.
Super-LIO Super-LIO* FAST-LIO2 FAST-LIO2* Faster-LIO iG-LIO
X86 3.98 3.12 0.84 1.32 0.60 1.66
ARM 1.71 1.37 0.39 0.66 0.40 0.43
To complement the average processing time, we also report a composite relative efficiency metric that jointly accounts for runtime and CPU usage:
$\eta=\frac{1}{\left(\tfrac{1}{N}\sum t_{i}\right)\left(\tfrac{1}{N}\sum u_{i}\right)}.$
Here, $t_{i}$ is the per-frame processing time (ms) and $u_{i}$ is the normalized CPU utilization (0-11). As summarized in Table IV, this metric consistently highlights the strong efficiency of Super-LIO across both platforms.
Figure 6: Module-level timing analysis of per-frame processing for the evaluated LIO frameworks. Each bar shows the time spent in the four core stages of the pipeline: undistortion, downsample, state update, and map update. Within each scene, the methods appear from left to right as: Super-LIO, Super-LIO*, Fast-LIO2, Fast-LIO2*, Faster-LIO, and iG-LIO.
Figure 7: State update time on representative sequences for OctVox, iVox [faster_lio], and iKD-Tree [fast_lio2] on X86 and ARM.
Figure 8: Map memory usage along the frame index on the NCLT 1 sequence.
Figure 6 shows the module-level timing across several representative scenes. Super-LIO maintains the best overall efficiency. The largest reduction comes from the state_update stage, highlighting the impact of OctVox and HKNN. Super-LIO also spends less time on undistortion due to parallel processing, whereas iG-LIO incurs additional overhead in downsample because of per-point covariance computations. Overall, the state_update stage dominates the total runtime across all frameworks.
To isolate the effect of the proposed map representation, Figure 7 replaces OctVox with iVox and iKD-Tree within the same Super-LIO pipeline and reports the average state update time on ten representative sequences for both X86 and ARM platforms. The improvements obtained with OctVox are clear under identical conditions. Finally, Figure 8 examines memory behavior on the longest sequence (nclt1, 1.8 hours). OctVox maintains the lowest and smoothest memory footprint without LRU, and with LRU enabled it stabilizes once reaching maximum capacity, while iVox shows minor memory variations.
Figure 9: Smoothed per-frame processing time on the NCLT 1 sequence.
Figure 10: CPU utilization over time on the NCLT 1 sequence.
Figures 9 and 10 examine the per-frame processing time and CPU usage on the NCLT 1 sequence. Super-LIO shows a clearly lighter computational load across the run. These plots provide a clearer view of the real-time behavior of all methods.
@@ IV-D Robustness Analysis
We also evaluated performance on both X86 and ARM platforms, confirming the cross-platform consistency of Super-LIO, as summarized in Tables II and III. In contrast, iG-LIO shows competitive real-time performance on X86 but runs slower on ARM and fails on all NTU VIRAL sequences.